OpenStudy (anonymous):

what is the inverse of the matrix??? [-8 5; 8 -3]

OpenStudy (zarkon):

\[\left[\begin{matrix}\frac{3}{16} & \frac{5}{16} \\ \frac{1}{2} & \frac{1}{2}\end{matrix}\right]\]